### v1

@@ 4.4 An Illustrative Example
We now apply the proposed framework to explain the agent’s behaviour in these two states, representing behaviour under unknown features by marginalising over the conditional steady-state distribution, as defined in Section 4.2. To do so, we first compute the steady-state distribution of the policy. Starting in State 1, the agent turns right, then left at State 2, and terminates thereafter. This yields a steady-state distribution of $0.5$ for each of the two states.
Using this distribution, we compute the agent’s behaviour under all subsets of observed features and use the resulting characteristic values to calculate Shapley values for each action. These are shown in Table 3.
Table 3: Characteristic values and Shapley values for each state-action pair $(s,a)$, computed by marginalising over the conditional steady-state distribution. Each state is denoted by its arrow direction (R for right, L for left) and the distance to the goal.
@@
R 00 00 00 $0.5$ $-0.25$ $-0.25$
The characteristic values and Shapley values are identical to those of Agent B in the earlier example: direction and distance contribute equally to the action taken in both visited states. This reflects our assumption—implicit in marginalising over the conditional steady-state distribution—that the agent uses all available feature information when acting. Accordingly, both features are treated as necessary for reproducing the agent’s policy, and responsibility is attributed proportionally.
This example demonstrates how defining the characteristic function by marginalising over the conditional steady-state distribution enables Shapley values to provide meaningful explanations of agent behaviour. [^2] See Section B.1 for an additional example illustrating how SVERL explains agent behaviour.
@@ 4.5 Explaining Behaviour in Continuous Action Spaces
A common policy representation in continuous control is a fixed-variance Gaussian, where actions are sampled from a distribution $\mathcal{N}(\mu(s),\sigma^{2})$, where $\mu:{\cal S}\rightarrow\mathbb{R}$ maps each state to the mean of the distribution. In this setting, a sound approach to explaining agent behaviour is to analyse how feature values determine the policy’s expected action $\mu(s)$. For example, how does a pendulum’s angular velocity influence the expected torque an agent applies in a pendulum balancing task?
This approach naturally leads to the question: how much does each individual feature contribute to the expected action $\mu(s)$? As before, we take a game-theoretic perspective, viewing $\mu(s)$ as the outcome of a coalitional game played by the state features. We identify Shapley values as the unique solution that satisfies four axioms defining a fair credit allocation.
@@
Central to the continuous policy game $({\cal F},\tilde{\mu}_{s})$ is the characteristic function $\tilde{\mu}_{s}$, which extends the mean function $\mu$ to situations where only a subset of state features is known. The precise construction of $\tilde{\mu}_{s}$ will be addressed later in this section; for now, we assume it is given and turn to our central question: what is the contribution of each feature value $s_{i}\in{\cal S}_{i}$ to the policy’s expected action $\mu(s)$? We denote this contribution by $\phi_{i}(\tilde{\mu}_{s})$. In other words, how can we fairly allocate—across the state features—the difference between the expected action when all features are known and when none are? Below, we define four more axioms that any attribution $\phi_{i}(\tilde{\mu}_{s})$, for all $i\in\mathcal{F}$, should satisfy to qualify as a fair credit allocation.
  • Axiom 1 (Efficiency). The difference in value between the expected action when all features are known, $\tilde{\mu}_{s}({\cal F})$, and the expected action when no features are known, $\tilde{\mu}_{s}(\emptyset)$, is fully distributed across the features:
    $\tilde{\mu}_{s}({\cal F})-\tilde{\mu}_{s}(\emptyset)=\sum_{i\in{\cal F}}\phi_{i}(\tilde{\mu}_{s})$
  • Axiom 2 (Linearity). For a composite policy that samples actions from a Gaussian whose mean $\alpha\,\tilde{\mu}(s)+(1-\alpha)\,\tilde{\nu}(s)$ is the weighted average of the mean functions $\tilde{\mu}(s)$ and $\tilde{\nu}(s)$, a feature’s contribution to the combined policy’s expected action is equal to the weighted sum of its contributions to each individual policy’s expected action:
    $\phi_{i}(\alpha\,\tilde{\mu}_{s}+(1-\alpha)\,\tilde{\nu}_{s})=\alpha\,\phi_{i}(\tilde{\mu}_{s})+(1-\alpha)\,\phi_{i}(\tilde{\nu}_{s})$
  • Axiom 3 (Symmetry). Two features are assigned the same contribution to the expected action if they make equal marginal contributions for all possible coalitions:
    $\phi_{i}(\tilde{\mu}_{s})=\phi_{j}(\tilde{\mu}_{s})\quad\text{if}\quad\tilde{\mu}_{s}(\mathcal{C}\cup\{i\})=\tilde{\mu}_{s}(\mathcal{C}\cup\{j\})\quad\forall\mathcal{C}\subseteq{\cal F}\setminus\{i,j\}$
  • Axiom 4 (Nullity). A feature is assigned zero contribution if it makes zero marginal contribution to the expected action for all possible coalitions:
    $\phi_{i}(\tilde{\mu}_{s})=0\quad\text{if}\quad\tilde{\mu}_{s}(\mathcal{C}\cup\{i\})=\tilde{\mu}_{s}(\mathcal{C})\quad\forall\mathcal{C}\subseteq{\cal F}\setminus\{i\}$
@@
$\tilde{\mu}_{s}(\mathcal{C})\equiv\omega(s_{\mathcal{C}})=\mathbb{E}[\mu(S)\mid S_{\mathcal{C}}=s_{\mathcal{C}}]=\sum_{s^{\prime}\in{\cal S}^{+}}p^{\pi}(s^{\prime}\mid s_{\mathcal{C}})\,\mu(s^{\prime}).$ (8)
@@
5 Explaining Performance
We now turn to explaining performance. In reinforcement learning, the performance of a policy $\pi$ is quantified by the state-value function $v^{\pi}:{\cal S}\rightarrow\mathbb{R}$, which gives the expected return from each state when following that policy:
@@ 5 Explaining Performance
Given that states are composed of feature values, it is natural to ask how individual features contribute to this expected return. For example, how does a navigation agent’s $x-$coordinate influence the expected return of its policy at a given location? In this section, we develop a theoretical framework for attributing expected return to the values of state features, offering a structured explanation of agent performance under policy $\pi$.
@@ 5.1 State Features Collaborate to Determine Expected Return
Definition 12 (Performance Game)
@@
The characteristic function $\tilde{v}^{\pi}_{s}$ extends the value function $v^{\pi}$ to settings where only a subset of feature values are accessible. As before, we postpone the precise construction of $\tilde{v}^{\pi}_{s}(\mathcal{C})$ until later. For now, we ask: what is the contribution of each feature value $s_{i}\in{\cal S}_{i}$ to the expected return $v^{\pi}(s)$? We denote this by $\phi_{i}(\tilde{v}^{\pi}_{s})$, define a set of fairness axioms that any such attribution should satisfy, and identify the Shapley values as the unique solution.
  • Axiom 1 (Efficiency). The difference in expected return between the policy with all feature values known and the policy with none known is fully distributed across the features:
    $\tilde{v}^{\pi}_{s}({\cal F})-\tilde{v}^{\pi}_{s}(\emptyset)=\sum_{i\in{\cal F}}\phi_{i}(\tilde{v}^{\pi}_{s})$
  • Axiom 2 (Linearity). For a composite policy $\alpha\,\pi+(1-\alpha)\,\omega$ that follows $\pi$ with probability $\alpha$ and $\omega$ with probability $1-\alpha$, a feature’s contribution to the expected return of the composite policy is equal to the same weighted combination of its contributions under each component policy:
    $\phi_{i}(\alpha\,\tilde{v}^{\pi}_{s}+(1-\alpha)\,\tilde{v}^{\omega}_{s})=\alpha\,\phi_{i}(\tilde{v}^{\pi}_{s})+(1-\alpha)\,\phi_{i}(\tilde{v}^{\omega}_{s})$
  • Axiom 3 (Symmetry). Two features are assigned the same contribution if they make equal marginal contributions to the expected return for all possible coalitions:
    $\phi_{i}(\tilde{v}^{\pi}_{s})=\phi_{j}(\tilde{v}^{\pi}_{s})\quad\text{if}\quad\tilde{v}^{\pi}_{s}(\mathcal{C}\cup\{i\})=\tilde{v}^{\pi}_{s}(\mathcal{C}\cup\{j\})\quad\forall\mathcal{C}\subseteq{\cal F}\setminus\{i,j\}$
  • Axiom 4 (Nullity). A feature is assigned zero contribution if it makes no marginal contribution to the expected return for all possible coalitions:
    $\phi_{i}(\tilde{v}^{\pi}_{s})=0\quad\text{if}\quad\tilde{v}^{\pi}_{s}(\mathcal{C}\cup\{i\})=\tilde{v}^{\pi}_{s}(\mathcal{C})\quad\forall\mathcal{C}\subseteq{\cal F}\setminus\{i\}$
@@
The Shapley values $\phi_{1}(\tilde{v}^{\pi}_{s}),\dots,\phi_{n}(\tilde{v}^{\pi}_{s})$ for the performance game $({\cal F},\tilde{v}^{\pi}_{s})$ are the unique allocation of the expected return $v^{\pi}(s)$ across the features of state $s$ that satisfies the four axioms of fair credit assignment: efficiency, linearity, symmetry, and nullity.
The Shapley values of the performance game provide a principled way to attribute expected return to individual state features, offering a concrete explanation of agent performance at a given state.
@@
We apply our framework to attribute the expected return at each state to its observed feature values. However, we must first make assumptions about how performance is affected by unknown features. We assume this corresponds to acting with partial information in the given state (at any decision step) and full observability elsewhere.
To specify how the agent acts with partial information, we reuse the two agents introduced earlier: Agent A follows the direction if visible and acts randomly otherwise. Agent B uses both direction and distance when available and acts randomly when neither is present. These assumptions define the policy under partial information and thus determine the expected return from each state for every feature subset. The resulting characteristic and Shapley values are shown in Tables 4 and 5.
@@
$(\text{R},10)$ $8$ $8$ $8$ $8$ 00 00
$(\text{L},2)$ $9$ $9$ $9$ $4$ $2.5$ $2.5$
Consider first the state $(\text{R},10)$. For both agents, neither feature contributes to performance. The reason is that turning left or right from the initial junction results in paths of equal length to the goal, yielding identical returns. In contrast, in our explanation of behaviour, the direction feature received credit for prompting the agent to turn right. This example demonstrates a key distinction between explaining behaviour and performance: a feature can influence behaviour without affecting performance.
In the second state, $(\text{L},2)$, the features do influence performance. For Agent A, all credit goes to the direction feature, reflecting its exclusive influence on the agent’s decision. For Agent B, both features contribute equally because either can be used to determine the correct action. These performance attributions mirror our behaviour analysis under the same assumptions, illustrating how the two forms of explanation can often align.
This example demonstrates how the proposed framework explains performance by attributing expected returns to individual state features. Crucially, the resulting explanations depend on the agent’s assumed behaviour under partial information. The following section outlines a general approach to addressing this challenge.
5.2 What Can Be Assumed About Performance When Some State Features Are Unknown?
We now consider how to define the characteristic function $\tilde{v}^{\pi}_{s}(\mathcal{C})$: the expected return from state $s$ when the agent’s policy has access only to the features in $\mathcal{C}$. Since the value function $v^{\pi}(s)$ assumes full observability, it does not specify this quantity, so we must make additional assumptions. If the agent’s expected return from state $s$ changes when certain feature values are unknown, this must reflect a shift in the agent’s behaviour at that state, assuming all else remains constant. Motivated by this, we define performance under unknown information as the expected return when the agent acts with partial information in state $s$ and follows the fully observed policy elsewhere. This definition isolates the contribution of features at state $s$ and allows us to express $\tilde{v}^{\pi}_{s}(\mathcal{C})$ using the policy characteristic function $\tilde{\pi}_{s}^{a}(\mathcal{C})$ introduced in Section 4.2.
Recall that $\tilde{\pi}_{s}^{a}(\mathcal{C})$ denotes the probability of selecting action $a$ in state $s$ when only the features in $\mathcal{C}$ are known:
@@ 5.2 What Can Be Assumed About Performance When Some State Features Are Unknown?
These probabilities form a complete action distribution at $s$ under partial information, which we can use to evaluate performance with unknown features.
Definition 14 (Performance Characteristic Function)
@@
This definition ensures that unknown features modify only behaviour at state $s$, isolating their contribution to performance. The agent behaves according to its original policy $\pi$ in all other states.
@@
Since we defined $\tilde{\pi}_{s}^{a}(\mathcal{C})$ as a marginal over the full state space, it may assign nonzero probability to unavailable actions in $s$. The distribution over ${\cal A}(s)$ must be renormalised in such cases.
@@
When only the features in $\mathcal{C}$ are available, the expected return $\tilde{v}^{\pi}_{s}(\mathcal{C})$ captures the outcome of following the best approximation to the original policy $\pi$ under this partial information. The difference between $\tilde{v}^{\pi}_{s}(\mathcal{C})$ and $v^{\pi}(s)$ thus quantifies how necessary the unknown features $\bar{\mathcal{C}}$ are for achieving the original expected return. Defining the characteristic function in this way ensures that the differences across feature subsets used to compute Shapley values capture the influence of the unknown features, making it a principled choice for the performance game. As before, this construction is general to any policy.
@@
Alternative definitions of agent behaviour under partial information could also be substituted into our proposed characteristic function. For example, the marginal variant of $\tilde{\pi}_{s}^{a}(\mathcal{C})$—which assumes feature independence—may be appropriate in some settings; see Section 4.3 for further discussion. Similarly, in continuous action spaces, the policy with partial information used in performance estimation can be defined as a Gaussian with mean $\mu_{s}(\mathcal{C})$, in place of the discrete policy based on $\tilde{\pi}_{s}^{a}(\mathcal{C})$; this substitution naturally extends the characteristic function $\tilde{v}^{\pi}_{s}(\mathcal{C})$ to continuous domains. We leave a detailed treatment of this extension to future work.
@@
We do not present a new illustrative example for the proposed characteristic function because its outcome is identical to the example in the previous subsection. In the road-sign domain, we noted how the policy characteristic $\tilde{\pi}_{s}^{a}(\mathcal{C})$ behaves exactly like Agent B under partial information. Since the performance characteristic $\tilde{v}^{\pi}_{s}(\mathcal{C})$ is defined using this same policy characteristic, it leads to the same values and attributions already shown for Agent B in Table 5. As such, the example already demonstrates how Shapley values enable faithful explanations of agent performance using this characteristic function. We provide an additional example illustrating how SVERL explains agent performance in Section B.2.
6 Explaining Value Estimation
We now turn to our third and final element of reinforcement learning that we will explain: value estimation. This element refers to how well an agent, or an observer, can estimate $v^{\pi}(s)$: the expected return of a policy $\pi$ from a given state. As before, our goal is to understand how the values of individual features contribute to this estimate.
At first glance, value estimation may seem indistinguishable from explaining performance: both involve analysing the expected return from a state. However, there is a crucial difference. When explaining performance, we asked how feature values influence the return an agent receives—by potentially altering the agent’s behaviour when some features are unknown. When explaining value estimation, by contrast, we ask how feature values influence the return an agent is estimated to receive—without altering how the agent behaves in that state.
To make this distinction precise, we introduce $\hat{v}^{\pi}:{\cal S}\rightarrow\mathbb{R}$ to denote expected return estimates from each state under policy $\pi$. When the estimate is accurate, $\hat{v}^{\pi}(s)$ equals $v^{\pi}(s)$, the true expected return from $s$. By changing our framing to treat $\hat{v}^{\pi}$ as a predictive function rather than an outcome, we present a theoretical analysis of explaining value estimation using the influence of feature values.

### v2

@@ 4.3 An Illustrative Example
We explain the agent’s behaviour in these two states by representing behaviour under unknown features by marginalising over the conditional steady-state distribution, as defined in Section 4.2. To do so, we first compute the steady-state distribution of the policy. Starting in state 1, the agent turns right, then left at state 2, and terminates thereafter. This yields a steady-state distribution probability of 0.50.50.5 for each of the two states. Using this distribution, we compute the agent’s behaviour under all subsets of observed features and use the resulting characteristic values to calculate Shapley values for each action, shown in Table 3. They reveal that direction and distance contribute equally to the action taken in both states. This result reflects the behaviour of marginalising over the conditional steady-state distribution: for this policy, either distance or direction is sufficient to identify the action taken, so they are treated as similarly informative and assigned symmetric credit.
@@
[…]
R 0 0 0 0.50.50.5 0.25-0.25- 0.25 0.25-0.25- 0.25
Table 3: Characteristic values and Shapley values for each state-action pair, computed by marginalising over the conditional steady-state distribution.
@@ 4.4 Explaining Behaviour in Continuous Action Spaces
A common policy representation in continuous control is a fixed-variance Gaussian, where actions are sampled from a distribution $\mathcal{N}(\mu(s),\sigma^{2})$, where $\mu:{\cal S}\rightarrow\mathbb{R}$ maps each state to the mean of the distribution. In this setting, a sound approach to explaining agent behaviour is to analyse how feature values determine the policy’s expected action $\mu(s)$. For example, how does a pendulum’s angular velocity influence the expected torque applied in a pendulum balancing task?
This approach naturally leads to the question: how much does each individual feature contribute to the expected action $\mu(s)$? As before, we take a game-theoretic perspective, viewing $\mu(s)$ as the outcome of a coalitional game played by the state features, with Shapley values as the unique solution that satisfies four axioms defining a fair credit allocation.
@@
Central to the continuous policy game $({\cal F},\tilde{\mu}_{s})$ is the characteristic function $\tilde{\mu}_{s}$, which extends the mean function $\mu$ to situations where only a subset of state features is known. The precise construction of $\tilde{\mu}_{s}$ will be addressed later in this section; for now, we assume it is given and turn to our central question: what is the contribution of each feature value $s_{i}\in{\cal S}_{i}$ to the policy’s expected action $\mu(s)$? We denote this contribution by $\phi_{i}(\tilde{\mu}_{s})$. In other words, how can we fairly allocate—across the state features—the difference between the expected action when all features are known and when none are? We again appeal to Shapley’s foundational axioms, restated here to clarify their role in this specific attribution setting. Below, we define four axioms that any attribution $\phi_{i}(\tilde{\mu}_{s})$, for all $i\in\mathcal{F}$, should satisfy to qualify as a fair credit allocation.
Axiom 20 (Efficiency)
The difference in value between the expected action when all features are known, $\tilde{\mu}_{s}({\cal F})$, and the expected action when no features are known, $\tilde{\mu}_{s}(\emptyset)$, is fully distributed across the features:
$\tilde{\mu}_{s}({\cal F})-\tilde{\mu}_{s}(\emptyset)=\sum_{i\in{\cal F}}\phi_{i}(\tilde{\mu}_{s}).$
Axiom 21 (Linearity)
For a composite policy that samples actions from a Gaussian whose mean $\alpha\,\tilde{\mu}(s)+(1-\alpha)\,\tilde{\nu}(s)$ is the weighted average of the mean functions $\tilde{\mu}(s)$ and $\tilde{\nu}(s)$ for two separate policies, a feature’s contribution to the combined policy’s expected action is equal to the weighted sum of its contributions to each individual policy’s expected action:
$\phi_{i}(\alpha\,\tilde{\mu}_{s}+(1-\alpha)\,\tilde{\nu}_{s})=\alpha\,\phi_{i}(\tilde{\mu}_{s})+(1-\alpha)\,\phi_{i}(\tilde{\nu}_{s}).$
Axiom 22 (Symmetry)
Two features are assigned the same contribution to the expected action if they make equal marginal contributions for all possible coalitions:
$\phi_{i}(\tilde{\mu}_{s})=\phi_{j}(\tilde{\mu}_{s})\quad\text{if}\quad\tilde{\mu}_{s}(\mathcal{C}\cup\{i\})=\tilde{\mu}_{s}(\mathcal{C}\cup\{j\})\quad\forall\mathcal{C}\subseteq{\cal F}\setminus\{i,j\}.$
Axiom 23 (Nullity)
A feature is assigned zero contribution if it makes zero marginal contribution to the expected action for all possible coalitions:
$\phi_{i}(\tilde{\mu}_{s})=0\quad\text{if}\quad\tilde{\mu}_{s}(\mathcal{C}\cup\{i\})=\tilde{\mu}_{s}(\mathcal{C})\quad\forall\mathcal{C}\subseteq{\cal F}\setminus\{i\}.$
@@
$\tilde{\mu}_{s}(\mathcal{C})\coloneqq\omega(s_{\mathcal{C}})=\mathbb{E}[\mu(S)\mid S_{\mathcal{C}}=s_{\mathcal{C}}]=\sum_{s^{\prime}\in{\cal S}^{+}}p^{\pi}(s^{\prime}\mid s_{\mathcal{C}})\,\mu(s^{\prime}).$ (8)
@@
5 Explaining Outcomes
Outcomes can refer to many measurable consequences of an agent’s behaviour, such as next states or immediate rewards. Here, we focus on expected return, a standard formalisation of outcome in reinforcement learning. Other forms of outcomes may be addressed in future work. The outcome of a policy $\pi$ can be quantified by the state-value function $v^{\pi}:{\cal S}\rightarrow\mathbb{R}$, which gives the expected return from each state when following that policy:
@@ 5 Explaining Outcomes
Given that states are composed of feature values, it is natural to ask how individual features contribute to this expected return. For example, how does a navigation agent’s $x-$coordinate influence the expected return of its policy at a given location? In this section, we develop a theoretical framework for attributing expected return to the values of state features, offering a structured explanation of agent outcomes under policy $\pi$.
@@ 5.1 State Features Collaborate to Determine Expected Return
Definition 28 (Outcome Game)
@@
The characteristic function $\tilde{v}^{\pi}_{s}$ extends the value function $v^{\pi}$ to settings where only a subset of feature values are accessible. As before, we postpone the precise construction of $\tilde{v}^{\pi}_{s}(\mathcal{C})$ until later. For now, we ask: what is the contribution of each feature value $s_{i}\in{\cal S}_{i}$ to the expected return $v^{\pi}(s)$? We denote this by $\phi_{i}(\tilde{v}^{\pi}_{s})$. We appeal to Shapley’s foundational axioms, restated here to clarify their role in this specific attribution setting, and identify the Shapley values as the unique solution.
Axiom 29 (Efficiency)
The difference in expected return between the policy with all features known and the policy with none known is fully distributed across the features:
$\tilde{v}^{\pi}_{s}({\cal F})-\tilde{v}^{\pi}_{s}(\emptyset)=\sum_{i\in{\cal F}}\phi_{i}(\tilde{v}^{\pi}_{s}).$
Axiom 30 (Linearity)
For two independent policies $\pi$ and $\omega$, a feature’s contribution to the weighted sum of their expected returns is equal to the corresponding weighted sum of the feature’s contributions under each policy:
$\phi_{i}(\alpha\,\tilde{v}^{\pi}_{s}+\beta\,\tilde{v}^{\omega}_{s})=\alpha\,\phi_{i}(\tilde{v}^{\pi}_{s})+\beta\,\phi_{i}(\tilde{v}^{\omega}_{s}).$
Axiom 31 (Symmetry)
Two features are assigned the same contribution if they make equal marginal contributions to the expected return for all possible coalitions:
$\phi_{i}(\tilde{v}^{\pi}_{s})=\phi_{j}(\tilde{v}^{\pi}_{s})\quad\text{if}\quad\tilde{v}^{\pi}_{s}(\mathcal{C}\cup\{i\})=\tilde{v}^{\pi}_{s}(\mathcal{C}\cup\{j\})\quad\forall\mathcal{C}\subseteq{\cal F}\setminus\{i,j\}.$
Axiom 32 (Nullity)
A feature is assigned zero contribution if it makes no marginal contribution to the expected return for all possible coalitions:
$\phi_{i}(\tilde{v}^{\pi}_{s})=0\quad\text{if}\quad\tilde{v}^{\pi}_{s}(\mathcal{C}\cup\{i\})=\tilde{v}^{\pi}_{s}(\mathcal{C})\quad\forall\mathcal{C}\subseteq{\cal F}\setminus\{i\}.$
@@
The Shapley values $\phi_{1}(\tilde{v}^{\pi}_{s}),\dots,\phi_{n}(\tilde{v}^{\pi}_{s})$ for the outcome game $({\cal F},\tilde{v}^{\pi}_{s})$ are the unique allocation of the expected return $v^{\pi}(s)$ across the features of state $s$ that satisfies the four axioms of fair credit assignment: efficiency, linearity, symmetry, and nullity.
The Shapley values of the outcome game provide a principled way to attribute expected return to individual state features, offering a concrete explanation of agent outcomes at a given state.
@@
We use the proposed framework to attribute the expected return from each state to its observed feature values. However, doing so requires an assumption about how unknown features affect expected return. To motivate this assumption, note that any change in expected return due to missing features must arise from a shift in the agent’s behaviour at that state, assuming behaviour and dynamics at all other states remain fixed. Based on this reasoning, we assume that the expected return with unknown features corresponds to the expected return from acting with partial information in the given state and full observability elsewhere. To specify how the agent acts with partial information, we reuse the two agents introduced earlier: Agent A follows the direction if visible and acts randomly otherwise. Agent B uses both direction and distance, turning right if it observes 10 miles or a right direction, left if it observes 2 miles or a left direction, and moving randomly otherwise. These assumptions define the policy under partial information and thus determine the expected return from each state for every feature subset. The resulting characteristic and Shapley values are shown in Tables 4 and 5.
[…]
@@
[…]
5.2 What Can Be Assumed About Expected Return When Some State Features Are Unknown?
We now consider how to define the characteristic function $\tilde{v}^{\pi}_{s}(\mathcal{C})$: the expected return from state $s$ when the agent’s policy has access only to the features in $\mathcal{C}$. Since the value function $v^{\pi}(s)$ assumes full observability, it does not specify this quantity, so we must make additional assumptions.
To motivate our definition, recall from the previous example that any change in expected return due to unknown features must reflect a shift in the agent’s behaviour at state $s$, assuming all else remains fixed. We therefore define expected return under unknown features as the expected return when the agent acts with partial information whenever it visits state $s$, while following its fully observed policy in all other states. This ensures that any change in return reflects only how the agent behaves differently at state $s$ under partial information, not how this might influence behaviour elsewhere in the environment. In this sense, the construction isolates the contribution of features at state $s$. It also enables us to express $\tilde{v}^{\pi}_{s}(\mathcal{C})$ using the discrete policy characteristic function $\tilde{\pi}_{s}^{a}(\mathcal{C})$ introduced in Section 4.2.
Recall that $\tilde{\pi}_{s}^{a}(\mathcal{C})$ defines the probability of selecting action $a$ in state $s$ when only the features in $\mathcal{C}$ are known:
@@ 5.2 What Can Be Assumed About Expected Return When Some State Features Are Unknown?
These probabilities form a complete action distribution at $s$ under partial information, which we can use to evaluate expected return with unknown features.
Definition 34 (Outcome Characteristic Function)
@@
This definition ensures that unknown features affect only how the agent behaves whenever it visits state $s$, while it follows its original policy $\pi$ in all other states. As a result, any change in expected return can be attributed specifically to the unknown features at state $s$, isolating their contribution.
@@
Since we defined $\tilde{\pi}_{s}^{a}(\mathcal{C})$ as a marginal over the full state space, it may assign nonzero probability to unavailable actions in $s$. In such cases, the distribution over ${\cal A}(s)$ should be renormalised with zero probability assigned to the unavailable actions.
@@
From Proposition 16, $\tilde{\pi}_{s}^{a}(\mathcal{C})$ is the best approximation to the original policy $\pi$ when only the features in $\mathcal{C}$ are available; the expected return $\tilde{v}^{\pi}_{s}(\mathcal{C})$ captures the outcome of following this approximation under the partial information. The difference between $\tilde{v}^{\pi}_{s}(\mathcal{C})$ and $v^{\pi}(s)$ thus quantifies how necessary the unknown features $\bar{\mathcal{C}}$ are for achieving the original expected return. Defining the characteristic function in this way ensures that the differences across feature subsets used to compute Shapley values capture the influence of the unknown features, making it a principled choice for the outcome game. As before, this construction is general to any policy.
@@
Alternative definitions of agent behaviour under partial information could also be substituted into our proposed characteristic function. For example, the marginal variant of $\tilde{\pi}_{s}^{a}(\mathcal{C})$—which assumes feature independence—may be appropriate in some settings; see Section 4.2 for further discussion. Similarly, in continuous action spaces, the policy with partial information can be defined as a Gaussian with mean $\mu_{s}(\mathcal{C})$, from Equation 8, in place of the discrete policy based on $\tilde{\pi}_{s}^{a}(\mathcal{C})$; this substitution naturally extends the characteristic function $\tilde{v}^{\pi}_{s}(\mathcal{C})$ to continuous domains. We leave a detailed treatment of this extension to future work.
@@
We do not present a new illustrative example for the proposed characteristic function because its outcome is identical to the example in the previous subsection. In the road-sign domain, the discrete policy characteristic $\tilde{\pi}_{s}^{a}(\mathcal{C})$ behaves exactly like Agent B under partial information. Since the outcome characteristic $\tilde{v}^{\pi}_{s}(\mathcal{C})$ is defined using this same policy characteristic, it leads to the same values and attributions already shown for Agent B in Table 5. As such, the example already demonstrates how Shapley values enable intuitive explanations of agent outcomes using this characteristic function. We provide an additional example illustrating how SVERL explains agent outcomes in Appendix B.
6 Explaining Predictions
Just as outcomes can refer to different consequences of agent behaviour, such as next states or rewards, predictions can target different estimates of these same quantities. In this section, we focus on predicting expected return, in line with our treatment of outcomes. Expected return can be formalised using either the state-value function $v^{\pi}(s)$ or the state-action value function $q^{\pi}(s,a)$; here, we focus on the former. Specifically, we consider how an agent or observer predicts $v^{\pi}(s)$, the expected return under policy $\pi$ from a given state. As before, we aim to understand how individual feature values contribute to this estimate. Analogous explanations for $q^{\pi}(s,a)$ and other prediction targets can be developed similarly.
At first glance, predictions may seem indistinguishable from explaining outcomes: both concern expected return under a given policy. However, the distinction lies in what role the feature values play. When explaining outcomes, we assess how features influence expected return through their effect on the agent’s behaviour. In contrast, when explaining predictions, we assess how features influence an external estimate of expected return, regardless of whether those features affect the agent’s behaviour.
To formalise this distinction, we introduce $\hat{v}^{\pi}:{\cal S}\rightarrow\mathbb{R}$, denoting an estimate of $v^{\pi}(s)$ made by the agent or an observer. When the estimate is accurate, $\hat{v}^{\pi}(s)=v^{\pi}(s)$, but our goal is not to evaluate its accuracy. Rather, by treating $\hat{v}^{\pi}$ as a predictive function, we isolate the epistemic question: how do feature values contribute to what we predict an agent to achieve, without assuming they shape what the agent actually does?